OpenStudy (anonymous):

Find the distance traveled in 15 seconds by an object traveling at a velocity of v(t) = 20 + 2cos(t) feet per second.

OpenStudy (anonymous):

\[x(t)=\int\limits_{}^{}v(t)dt\]

OpenStudy (anonymous):

\[x(t)=\int\limits_{0}^{15}20dt+\int\limits_{0}^{15}2\cos(t)dt\]

OpenStudy (anonymous):

solve the integral

OpenStudy (anonymous):

When i solve for the integral I am getting an error screen. What is the answer that you get?

OpenStudy (anonymous):

integrate using rules of integration instead of a calculator...

OpenStudy (anonymous):

From Mathematica 9: \[\int\limits_0^{15} 20 \, dt+\int\limits_0^{15} 2 \cos (t) \, dt=300+2 \sin (15)=301.301 \]